• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 12:32:58 on August 21, 2009

    Entry number 285132

    keyword=Right Spectrometer Dipole work

    The dipole seems to be working correctly. The maximum output voltage is set at 3 volts. I replaced the socket and opamp for U2 on the Regulator board. While working on the dipole power supply, Javier and I found the output pin was making intermittent contact with the socket and pin 8 had been folded up under the IC due to not being inserted carefully. The small current offset will be different than it has been by an amp or two. We also found the serial cable to the Octagon controller had a poor connection at a jumper cable inside the supply. The jumper was clearly a temporary fix that never was improved. We made a different jumper and ordered parts for a more robust fix. While removing a wire from one of the DB25 breakout boards in the supply the screw terminal was damaged. Most of the terminals seemed to be over-tightened. Two more screw terminals were damaged when we tried to loosen the screw. A temporary replacement breakout board was installed and the correct ones are on order.